determine the pH 0.270 M NH4NO3 and 0.101 M HCN

Using an ice table
Answer #1

NH4 + H₂O ² NH₃ + H₂O 1 0.27M - c -a + 4x2 0.27 m -x tn tn Ra= 5.88×10 ka = [NHI] [Hot] Inho] Sorrx100x² (0.27-) 60.27 >> 0.2